OpenStudy (anonymous):

Match each equation with its y-intercep 6x-5y=30 A.-6 B.-5 C.5 D.6

OpenStudy (anonymous):

a y-intercept is a point on the graph where x is zero.

OpenStudy (anonymous):

So what would be the answer ?

OpenStudy (anonymous):

A.-6

OpenStudy (anonymous):

this is the correct answer this time as 6*0-5*6=30 30/5=6 -5*-6=30